OpenStudy (anonymous):

x+3/x-4 + x/4-x add and simplify if possible

OpenStudy (anonymous):



OpenStudy (anonymous):

i dont understand what im supposed to do with these kind of problems and the way its set up

hero (hero):

Hint: \(\large\frac{x+3}{x - 4} = -\frac{x+3}{4-x}\)

hero (hero):

So: Replacing that, you get \(-\large\frac{x+3}{4-x} + \frac{x}{4-x} \) now just add

hero (hero):

carefully

hero (hero):

I'm basically showing you the stuff books won't show you

OpenStudy (anonymous):

3 ----- 4 - x ?

hero (hero):

Almost, you made a slight mistake. It's actually \(-\large\frac{3}{4-x} \)

OpenStudy (anonymous):

its a multiple choice and thats not an option but 3 over (4 - x) is

hero (hero):

What are the other options?

hero (hero):

The answer is \(-\frac{3}{4-x}\) but even the book can make mistakes

OpenStudy (anonymous):

(2x + 3) 2x+ 3 ------- and ------- x - 4 4- x

OpenStudy (anonymous):

3 ---- 4 - x

OpenStudy (anonymous):

those are the three options

hero (hero):

lol, the book made a mistake

hero (hero):

I don't know what to tell you

OpenStudy (anonymous):

okay.. thanks

hero (hero):

actually it is \(\frac{3}{x-4}\)

hero (hero):

Somehow I goofed up on the negative

hero (hero):

Oh, I see

hero (hero):

What I posted was also correct

hero (hero):

\(-\frac{3}{4-x} = \frac{3}{x-4}\)